OpenStudy (anonymous):

What is the sum of a 22-term arithmetic sequence where the first term is 54 and the last term is -30? (1 points) 264 288 312 336

hartnn (hartnn):

use this direct formula \(\Large S_n = \dfrac{n}{2}(a_1+a_n)\)

hartnn (hartnn):

where n = number of terms = 22 a1 = 1st term = 54 an = last term = -30 just plug in!

OpenStudy (anonymous):

Ok

hartnn (hartnn):

calculate and tell me what you get, i'll verify the answer for you :)

OpenStudy (anonymous):

So its \[ \frac{ 22 }{ 2 } \left( 54+ -30\right)\]

hartnn (hartnn):

correct, go on!

hartnn (hartnn):

54-30 = 24

OpenStudy (anonymous):

264

hartnn (hartnn):

\(\huge \checkmark \)

OpenStudy (anonymous):

OMG YAY THANK YOU SO MUCH

hartnn (hartnn):

welcome ^_^